OpenStudy (anonymous):

State the xintercept and yintercept for the graph of each equation. Then, graph the equation using the xand yintercepts. 1. 2x + 3y = 12 2. y = 2x - 4

OpenStudy (hockeychick23):

ok so to solve the x intercept we set y equal to zero. So: 2x + 3(0) = 12 2x=12 x intercept=6 Now we do the same for y. 2(0)+3y=12 y intercept=4

OpenStudy (hockeychick23):

we do the same thing for 2. 0 = 2x - 4 4= 2x x intercept=2 y=2(0)-4 y intercept= -4